# Inferring Foresightedness in Dynamic Noncooperative Games

### v2

@@ VI-E Monte Carlo Results - Intersection
As previously discussed, Figure 2 shows a representative trial of our method and the baseline with the real-world data from the InD dataset [inDdataset]. It shows the initial positions, the perceived and inferred goals, the observations and recovered trajectories up to a time $t$ = 15, and the planned trajectories for all four players at $t$ = 15 for the baseline and our method. Our visualization shows a much closer recovery of the ground-truth goal position than the baseline. In addition, we recover a visually similar trajectory when compared to the observations.
@@
Figure 5 shows the mean trajectory error as a function of the observation noise standard deviation for the previously described methods. Our method consistently shows a significant improvement on the baseline in both the fully observable and partially observable cases, leading to more accurate trajectory recovery. However, both methods seem to be relatively insensitive to increases in noise, as they create only a small increase in error. Furthermore, unlike the synthetic data case, both methods are also relatively consistent in terms of error standard deviation.
Since we have introduced only modest levels of noise that are broadly consistent with real-world sensors, we attribute the noise-insensitivity of Figure 5 to model mismatch. Model mismatch refers to the differences between actual human decision-making processes and our recreation of human decision-making. Two dominant sources of mismatch are likely to be in the information structure of our recreation and the expressiveness of our cost structures. In this paper, we assume an open-loop information structure instead of the more human closed-loop structure for tractability. While more complex cost structures may close the gap between our model and human decision-making, they are unlikely to be more significant than the information structure mismatch.
VI-F Receding Horizon
Figure 6: Ratio of baseline error over our method’s error when estimating players’ goal positions in a receding horizon setup. Results were obtained from an experiment with noise standard deviation 0.002 $\mathrm{m}$.
In practice, we envision solving 1 in a receding horizon fashion. Figure 2 shows a sample implementation of this type of solution. Our results from a Monte Carlo study for a single noise level (standard deviation 0.002 $\mathrm{m}$) and 50 trials demonstrate a largely improved performance over the baseline in this mode of operation.
Figure 6 shows the error in estimating the goal position at each time step for each player. Specifically, the vertical axis shows the ratio between the baseline’s error and ours, $\frac{\|\theta_{\text{goal, our method}}-\theta_{\textrm{goal, ground-truth}}\|^{2}_{2}}{\|\theta_{\text{goal, baseline}}-\theta_{\textrm{goal, ground-truth}}\|^{2}_{2}}$ (ratios greater than 1 show that our method improves on the baseline). As $\theta_{\textrm{goal, ground-truth}}$ is unknown, we use real-world data of each agent’s last known position as a proxy. Notably, our method generally exhibits lower goal position error and higher noise robustness across all players and across most planning steps—this promising result indicates that our method yields more accurate predictions for receding horizon planning models than the baseline, which ignores foresightedness.
However, notice that at time step $15$, the baseline sees a large drop in error while our method experiences a slight increase. This behavior corresponds to the time when the agent waiting for the pedestrians to cross finally starts moving again. The direction and magnitude of agents’ motion is highly indicative of their desired goal positions.
@@ VII Conclusions
This paper formulates a noncooperative game that models agents as potentially foresighted by associating each agent with an unknown discount factor. Our approach rewrites the equilibrium conditions of the resulting game as an MiCP and leverages the directional differentiability of equilibrium solutions with respect to unknown parameters. This construction allows us to solve an inverse game problem and identify hidden parameters—including agents’ unknown discount factors—via a gradient descent procedure. By evaluating our method with noisy partial and full state observations on a simulated crosswalk environment and a real-world intersection, we demonstrate superior performance compared to an existing state-of-the-art method.
[…]

### v3

@@ VI-F Monte Carlo Results - InD Intersection
As previously discussed, Figure 2 shows a representative trial of our method and the baseline with the real-world data from the InD dataset [48]. It shows the initial positions, the perceived and inferred goals, the observations and recovered trajectories up to a time $t$ = 15, and the planned trajectories for all four players at $t$ = 15 for the baseline and our method. Our visualization shows a closer recovery of the ground-truth goal position than the baseline. In addition, we recover a trajectory closer to the observations in L2 distance than baseline.
@@
Figure 6 shows the mean trajectory error as a function of the observation noise standard deviation for the previously described methods. On average, our method consistently shows a significant improvement on the baseline in both the fully observable and partially observable cases. However, both methods seem to be relatively insensitive to increases in noise, as they create only a small increase in average error. Furthermore, unlike the synthetic data case, both methods are also relatively consistent in terms of error standard deviation.
Note that in both the full and partial state experiments, the noiseless-observation trajectory error is non-zero. We attribute these errors to model mismatch, referring to the differences between actual human decision-making processes and our recreation of human decision-making. Two likely dominant sources of mismatch are our use of an open-loop information structure—chosen for tractability over a more human closed-loop form—and the expressiveness of our cost structures. More complex cost structures, such as those encoded by neural networks, may also close the gap.
(a) Only position measurements
(b) Full state measurements
Figure 6: Bootstrapped estimates (solid lines) and standard errors for the average trajectory prediction error for the baseline method and our method for the InD intersection experiment. On average, our method produces trajectories with less error in settings with both (a) partial state observations and (b) full state observations.
VI-G Monte Carlo Results - Waymax
Finally, we show how our method can be integrated into a receding horizon motion planner, and present several metrics related to safety and efficiency. In Figure 7, we show that a foresighted game-theoretic receding horizon planner maintains a distance of over $6\text{\,}\mathrm{m}$ from all cars at all times across all noise realizations, exhibiting a high level of safety. Meanwhile, the ego agent is reliably able to minimize the distance to its goal across all trials. These results demonstrate that our foresighted formulation yields safe and efficient behavior when employed in receding horizon robotics schemes. A representative trial of this simulation is shown in Figure 3.
@@ VII Conclusions
This paper formulates a noncooperative game that models potentially foresighted agents by associating each one with an unknown discount factor. Our approach rewrites the equilibrium conditions of the resulting game as an MiCP to leverage the directional differentiability of its equilibrium solutions with respect to unknown parameters. This construction allows us to solve an inverse game problem and identify hidden parameters—including agents’ unknown discount factors—via a gradient descent procedure. By evaluating our method with noisy partial and full state observations on a simulated crosswalk environment and a real-world intersection and road, we demonstrate superior performance compared to an existing state-of-the-art method.
This work models foresighted behavior through the use of discount factors. In this vein, we find several avenues for further research interesting. First, because neural networks are universal function approximators, they can be used to approximate each player’s cost function. This could potentially eliminate model mismatch and show greater improvement on our method. Second, humans often make decisions under the assumption they will have more information in the future, i.e., within feedback information structures. As noted in Section VI-F, future work may extend our open-loop formulation to better model this behavior, with [19] offering a clear starting point for this direction. Third, humans may adapt how far ahead they plan based on context. On clear highways, they may focus on the near term, but in complex situations like merging or heavy traffic, they consider longer horizons. Future work may explore discounting schemes that vary with state and time.
Figure 7: Several metrics computed on 5050 trajectories generated from the Waymax Monte Carlo simulation with an observation noise level of $0.1\text{\,}{\mathrm{m}}^{2}$. In particular, we focus on the distance between the ego robot and the other most important drivers in the simulation, as well as how far the ego robot is from their goal.